If a,b,c are three consecutive integers , Prove that (a-i)(a+i)(c+i)(c-i)=b^4 +1

Grade:11

1 Answers

Aditya Gupta
5 years ago
LHS= (1+a^2)(1+c^2)=1+a^2+c^2+a^2c^2
But c=b+1 and a=b-1
1+2(b^2+1)+(b^2-1)^2
= 1+2(b^2+1)+ b^4+1-2b^2
= b^4+4 and not b^4+1